# A finite volume Simo-Reissner beam method for moored floating body dynamics

### v1

@@ 1 Introduction
The importance of mooring lines extends beyond mere attachment and station-keeping; they fundamentally contribute to the safety and operational viability of floating offshore structures, especially in environments where uncontrolled movement could lead to catastrophic consequences, such as complete structural failure [1]. An accurate model of mooring lines is not only essential for understanding the hydrodynamic response amplitude of floating offshore structures but is also critical for assessing long-term fatigue and ensuring the longevity of both the mooring system and the floating structure it supports.
@@
Mooring line analysis models are generally categorised into static, quasi-static, and dynamic types. Static models [2, 3] assume that the system is in static equilibrium, accounting only for constant forces such as steady wind or current loads while neglecting any time-dependent or dynamic effects. This simplification allows for basic assessments of mooring line tension and position [4] under calm or steady-state conditions but may overlook critical forces in more variable environments. Quasi-static models, on the other hand, enhance the realism of static analyses by incorporating slowly varying environmental forces and accounting for the mooring line’s stiffness response to changing loads [5, 6]. However, they still neglect high-frequency dynamics associated with wave action. The quasi-static approach offers a balance between simplicity and accuracy [7]. It is well-suited for conditions dominated by low-frequency forces. In contrast, full dynamic modelling is more appropriate for scenarios where wave-induced motions play a significant role. When the floating structures experience large amplitude motions, the dynamic behaviours of the mooring system become non-negligible [8]. Thus, quasi-static models will underestimate the tension on the lines [9, 10]. Incorporating dynamic effects into the equations of motion for mooring lines allows dynamic analysis models to predict mooring loads and the resulting responses of floating structures with greater accuracy compared to static and quasi-static models. Mooring Design and Dynamics (MDD) [11] which solves the dynamic equilibrium equations under environmental loads, SEAWAY [12], which is a frequency-domain model, OrcaFlex [13] which is a non-linear framework for time-domain analysis, are some of the recent dynamic mooring models. Ferri and Palm [14] developed a finite element model (named Moody) with high-order polynomial basis functions and spatial discretisation using the local discontinuous Galerkin method. Following this higher-order approach, Moody was developed with the aim of resolving snap loads, and it incorporates a range of hydrodynamic forces within the model [15]. The bending stiffness has later been added to the model to investigate the bending stiffness on snap-loads [16]. Hall et al. [17] developed an open-source dynamic model called MoorDyn based on lumped mass approach [18], which discretises the mooring line into a finite number of mass nodes connected with mass-less springs. While the initial version of MoorDyn neglects bending stiffness, this feature was later added in the model [19]. In all the aforementioned references, the consistent conclusion is that incorporating bending stiffness into mooring cable analysis diversifies the usability of the numerical simulation to low-tension and slack conditions, avoiding any ill-conditioning of the system characteristics.
@@
Recently, some researchers have focused on coupling dynamic mooring analysis models with computational fluid dynamics (CFD) simulations to enhance the accuracy of mooring system behaviour predictions under complex fluid-structure interactions. This approach allows for the detailed simulation of forces from waves and currents on mooring lines, improving the understanding of both static and dynamic responses in challenging offshore environments [20]. CFD models can utilise either mesh-based or mesh-free approaches. Mesh-free approaches are primarily suitable for free-surface wave structure interaction problems due to their speed-up when using graphical computing units and their ability to capture interfaces without any additional processing [9, 21]. Dominguez et al. [22] was the first to couple DualSPHysics [23], which is an open-source parallel mesh-free CFD solver based on smoothed particle hydrodynamics (SPH) with MoorDyn. Later Liu and Wang [24] extended this coupling to study different submerged box-type breakwaters. Although these mesh-free methods are popular among researchers, they can become highly complex regarding particle refinement and turbulence modelling [25, 26]. Islam et al. [27] studied the wave radiation due to heave, surge, and pitch by a box-type floating structure using OpenFOAM [28], a widely-used CFD toolbox. Haider et al. [29] carried out a fully coupled fluid-structure interaction analysis for a floating offshore wind turbine, employing MoorDyn to model the mooring system and OpenFOAM as the CFD platform. Wei et al. [30] performed a coupled simulation of a very large floating structure (VLFS) moored with catenary mooring lines using OpenFOAM and MBDyn [31]. Palm et al. [32] conducted a coupled CFD-mooring analysis of a wave energy converter (WEC), coupling OpenFOAM with Moody to simulate the behaviour of the mooring lines. Similarly, Chen and Hall [9] integrated MoorDyn with OpenFOAM to simulate a moored floating box. More recently, they extended their work to model multiple floating objects [33]. Beyond the studies mentioned, additional research has explored coupling mooring line models with various CFD software [34, 35, 36, 37]. However, OpenFOAM’s customizability makes it a preferred choice among researchers as a CFD solver. Furthermore, the reviewed studies highlight OpenFOAM’s capability to accurately simulate the hydrodynamic response of floating rigid bodies.
@@
The literature on moorings highlights the importance of incorporating bending stiffness into numerical simulations for accurate analysis. For instance, commercial software such as OrcaFlex [38], which employs a lumped mass approach, explicitly calculates bending moments based on the bending angle at the intersections of the lumped nodes in the mooring line model. Rather than using lumped mass or finite element-based approaches, one can consider the mooring line as a beam or rod model undergoing deformations due to its slender shape and flexibility [39]. In the existing literature, beam models range from linear formulations like the Euler-Bernoulli model, which neglects shear deformation, to the geometrically exact Simo-Reissner beam model. The latter are non-linear and account for large deformations and rotations, offering a more comprehensive representation of beam behaviour [40, 41, 42]. While numerous finite element formulations [43, 44, 45, 46] have been introduced in the literature for the aforementioned geometrically exact Simo-Reissner beam theory, Tukovic et al. [47] was to first to propose a finite volume formulation of geometrically exact Simo-Reissner beams. More recently, Bali et al. [48] expanded the finite volume formulation for shear-deformable geometrically exact beams and conducted a comparative analysis with finite element-based classical benchmark cases using OpenFOAM. Additionally, they extended this formulation to handle beam-to-beam contact interactions [49]. The primary motivation for adapting the Simo–Reissner formulation to finite volume discretisation within OpenFOAM is to unify numerical methods for solids and fluids, allowing complex multi-physics problems, like fluid-solid interactions, to be addressed within a unified framework. Furthermore, it has been noted [48] that, unlike the finite element approach, the finite volume formulation for beams is not subject to the shear-locking phenomenon.
@@
In the current work, the finite volume Simo Reissner beam formulation introduced by Bali et al. [48] is extended by fully integrating it with a multiphase CFD solver and a six degree of freedom motion solver. The resulting framework captures large nonlinear deformations in mooring lines, including bending, shear and torsion, together with the fluid’s free surface dynamics in a single finite volume context. Building on these previous efforts and addressing the need for a unified solver, this novel integration is fully implemented within a finite volume framework, resulting in a non-linear, coupled dynamic model for moored floating structures that captures the complex interactions between structural dynamics and environmental forces. The model also incorporates seabed contact and seabed friction to more accurately represent scenarios where mooring lines rest on or interact with the seabed. The coupled model is validated and verified by comparing its results with both experimental measurements [50] and numerical simulations from MoorDyn [9] for a floating box moored with catenary lines under regular waves, obtained from an introductory test campaign for wave energy converters. The rest of the paper is organised as follows. Section 2 provides the details of the mathematical model of the moored floating body problem and the coupling strategy. Section 3 elaborates on the benchmarking of the proposed algorithm. Finally, some conclusions are drawn in Section 4.
@@ 2 Mathematical Model and Numerical Methods
The proposed approach utilises OpenFOAM [28] as the main framework for CFD analysis, while the beam solver, also implemented within the OpenFOAM framework, is used for mooring line dynamics modelling. Consequently, the equation of motion for the floating rigid body and its interaction with mooring lines must be coupled with the CFD model. This section elaborates on the mathematical models that have been used for each component of a moored floating body problem.
@@ 2.1 Free surface flows
The volume of fluid (VOF) formulation based on the surface capturing method with Reynolds Averaging Navier-Stokes (RANS) equations is employed here for solving a two-phase air-water system. The continuity and the conservation of momentum equation are given by:
@@
Physically, Eqn. 5 represents the conservation of mass for one of the phases in a multiphase flow, ensuring that the volume fraction evolves correctly over time. The Multidimensional Universal Limiter with Explicit Solution (MULES) algorithm [51] is applied here to reduce interface diffusion. This method introduces an artificial compression (or anti-diffusion) term $\nabla\cdot[\mathbf{U_{r}}\alpha(1-\alpha)]$ on the left side of Eqn. 5. In this term, $\mathbf{U_{r}}$ represents the compression velocity. The final expression for the volume fraction advection equation then becomes:
@@
Further detail on MULES formulation can be found in Damián and Nigro [51].
@@ 2.2 Mooring model
$\frac{\partial\mathbf{m}}{\partial s}+\mathbf{r}^{\prime}\times\mathbf{n}+\bm{\tau}_{\text{ext}}=\frac{\partial}{\partial t}\left(\mathbf{I}\bm{\omega}\right)$ (8)
where in Eqn. 7, $\mathbf{n}(s,t)$ represents the internal force resultant vector at position $s$ and time $t$, $\mathbf{f}_{ext}(s,t)$ is the external force per unit length acting on the beam at position $s$ and time $t$, $\rho_{b}$ is the density of the beam, and $A$ is the cross-sectional area of the beam. Consequently $\rho_{b}A$ is the mass per unit length of the beam, $\frac{\partial^{2}\mathbf{r}}{\partial t^{2}}$ is the linear acceleration of the point on the centre line at position $s$. In the moment equation (Eqn. 8), $\mathbf{m}$ is the internal moment vector, $\mathbf{r}^{\prime}$ is the tangent vector along the beam’s centre line, $\bm{\tau}_{\text{ext}}(s,t)$ represents the external torques, $\mathbf{I}$ is the moment of inertia tensor of the beam’s cross-section and $\bm{\omega}$ is the angular velocity vector. Equation 7 represents the conservation of linear momentum, where the net internal and external forces acting on a beam segment result in its linear acceleration. Similarly, Eqn. 8 represents the conservation of angular momentum, accounting for the net internal moments and external torques acting on the segment.
@@
The drag force $\mathbf{f}_{d}$ based on Morrison’s equation [52], which describes the total hydrodynamic forces acting on a structure submerged in or interacting with fluid flow, comprises of normal and tangent components expressed as [32]
@@
$\epsilon=\left|\left|\frac{\partial\mathbf{r}}{\partial s}\right|\right|-1$ (11)
$\mathbf{V}_{r,t}$ and $\mathbf{V}_{r,n}$ are the tangential and normal components of the relative velocity between the beam and the surrounding fluid, i.e. $\mathbf{V}_{r}=\mathbf{V}-\mathbf{V}_{fluid}$. In this study, a quiescent fluid is assumed around the mooring line, i.e. $\mathbf{V}_{fluid}=0$. Similar to the drag force, the added mass force is one of the components in the Morison equation [52]. The added mass force arises due to the fact that when a body moves through a fluid, it must push aside or accelerate the surrounding fluid. This extra inertia, which the body feels due to the fluid’s resistance, is called added mass. This term increases the effective mass of the beam because it includes both the actual mass of the beam and the mass of the fluid that the beam displaces during motion. Following the similar notation in Eqn. 10, the added mass force can be expressed as
@@
The total buoyancy force acting on the beam proportional to the displaced fluid can be expressed as [32]
@@
where $\mathbf{A}_{w}$, $\mathbf{A}_{c}$, and $\mathbf{A}_{e}$ are $6\times 6$ coefficient matrices; $\mathbf{X}_{w}$, $\mathbf{X}_{c}$, and $\mathbf{X}_{e}$ are $6\times 1$ vectors of unknown displacements and rotations, and $\mathbf{B}_{c}$ is the corresponding $6\times 1$ source vector. This form is assembled for each beam cell, leading to a global system of $6\times N$ coupled equations. The resulting global matrix is sparse and block tri-diagonal, and the non-linear system is solved using the Newton–Raphson method. The detailed formulation of the Simo-Reissner beam theory, alongside their discretisation, can be found in Bali et al. [48].

### v2

@@ 1 Introduction
The importance of mooring lines extends beyond mere attachment and station-keeping; they fundamentally contribute to the safety and operational viability of floating offshore structures, especially in environments where uncontrolled movement could lead to catastrophic consequences, such as complete structural failure [yang_investigation_2021]. An accurate model of mooring lines is not only essential for understanding the hydrodynamic response amplitude of floating offshore structures but is also critical for assessing long-term fatigue and ensuring the longevity of both the mooring system and the floating structure it supports.
@@
Mooring line analysis models are generally categorised into static, quasi-static, and dynamic types. Static models [oppenheim_static_1982, smith_statics_2001] assume that the system is in static equilibrium, accounting only for constant forces such as steady wind or current loads while neglecting any time-dependent or dynamic effects. This simplification allows for basic assessments of mooring line tension and position [karnoski_validation_1988] under calm or steady-state conditions but may overlook critical forces in more variable environments. Quasi-static models, on the other hand, enhance the realism of static analyses by incorporating slowly varying environmental forces and accounting for the mooring line’s stiffness response to changing loads [dunbar_development_2015, cheng_numerical_2019]. However, they still neglect high-frequency dynamics associated with wave action. The quasi-static approach offers a balance between simplicity and accuracy [liu_establishing_2017]. It is well-suited for conditions dominated by low-frequency forces. In contrast, full dynamic modelling is more appropriate for scenarios where wave-induced motions play a significant role. When the floating structures experience large amplitude motions, the dynamic behaviours of the mooring system become non-negligible [masciola_implementation_2013]. Thus, quasi-static models will underestimate the tension on the lines [chen_cfd_2022, hall_validation_2015]. Incorporating dynamic effects into the equations of motion for mooring lines allows dynamic analysis models to predict mooring loads and the resulting responses of floating structures with greater accuracy compared to static and quasi-static models. Recent dynamic mooring models include Mooring Design and Dynamics [dewey_mooring_1999], which addresses dynamic equilibrium equations under environmental loads; SEAWAY [jmj_journee_ljm_adegeest_theoretical_2003], a frequency-domain model; and OrcaFlex [randolph_non-linear_2010], a non-linear framework for time-domain analysis. Ferri and Palm [ferri_implementation_2015] developed a finite element model (named Moody) with high-order polynomial basis functions and spatial discretisation using the local discontinuous Galerkin method. Following this higher-order approach, Moody was developed with the aim of resolving snap loads, and it incorporates a range of hydrodynamic forces within the model [davidson_mathematical_2017]. The bending stiffness has later been added to the model to investigate the bending stiffness on snap-loads [palm_influence_2020]. Hall et al. [hall_moordyn_2015] developed an open-source dynamic model called MoorDyn based on a lumped mass approach [kreuzer_mooring_2002], which discretises the mooring line into a finite number of mass nodes connected with mass-less springs. While the initial version of MoorDyn neglects bending stiffness, this feature was later added in the model [hall_implementation_2021]. In all the aforementioned references, the consistent conclusion is that incorporating bending stiffness into mooring cable analysis diversifies the usability of the numerical simulation to low-tension and slack conditions, avoiding any ill-conditioning of the system characteristics.
@@
Recently, some researchers have focused on coupling dynamic mooring analysis models with computational fluid dynamics (CFD) simulations to enhance the accuracy of mooring system behaviour predictions under complex fluid-structure interactions. This approach allows for the detailed simulation of forces from waves and currents on mooring lines, improving the understanding of both static and dynamic responses in challenging offshore environments [darling_role_2024]. CFD models can utilise either mesh-based or mesh-free approaches. Mesh-free approaches are primarily suitable for free-surface wave structure interaction problems due to their speed-up when using graphical computing units and their ability to capture interfaces without any additional processing [chen_cfd_2022, lind_review_2020]. Dominguez et al. [dominguez_sph_2019] was the first to couple DualSPHysics [crespo_dualsphysics_2015], which is an open-source parallel mesh-free CFD solver based on smoothed particle hydrodynamics (SPH) with MoorDyn. Later, Liu and Wang [liu_numerical_2020] extended this coupling to study different submerged box-type breakwaters. Although these mesh-free methods are popular among researchers, they can become highly complex regarding particle refinement and turbulence modelling [macia_truncated_2021, vacondio_grand_2021]. Islam et al. [islam_openfoam_2019] studied the wave radiation due to heave, surge, and pitch by a box-type floating structure using OpenFOAM [weller_tensorial_1998], a widely-used CFD toolbox. Haider et al. [haider_comprehensive_2024] carried out a fully coupled fluid-structure interaction analysis for a floating offshore wind turbine, employing MoorDyn to model the mooring system and OpenFOAM as the CFD platform. Wei et al. [wei_coupled_2024] performed a coupled simulation of a very large floating structure (VLFS) moored with catenary mooring lines using OpenFOAM and MBDyn [masarati_efficient_2014]. Palm et al. [palm_coupled_2016] conducted a coupled CFD-mooring analysis of a wave energy converter (WEC), coupling OpenFOAM with Moody to simulate the behaviour of the mooring lines. Similarly, Chen and Hall [chen_cfd_2022] integrated MoorDyn with OpenFOAM to simulate a moored floating box. More recently, they extended their work to model multiple floating objects [CHEN2024117697]. Beyond the studies mentioned, additional research has explored coupling mooring line models with various CFD software [qiao_dynamic_2024, huang_new_2022, martin_numerical_2021, chen_coupled_2018]. However, OpenFOAM’s customizability makes it a preferred choice among researchers as a CFD solver. Furthermore, the reviewed studies highlight OpenFOAM’s capability to accurately simulate the hydrodynamic response of floating rigid bodies.
@@
The literature on moorings highlights the importance of incorporating bending stiffness into numerical simulations for accurate analysis. For instance, commercial software such as OrcaFlex [orcina_orcaflexdocumentation_2020], which employs a lumped mass approach, explicitly calculates bending moments based on the bending angle at the intersections of the lumped nodes in the mooring line model. Rather than using lumped mass or finite element-based approaches, one can consider the mooring line as a beam or rod model undergoing deformations due to its slender shape and flexibility [cottanceau_finite_2018]. In the existing literature, beam models range from linear formulations like the Euler-Bernoulli model, which neglects shear deformation, to the geometrically exact Simo-Reissner beam model. The latter are non-linear and account for large deformations and rotations, offering a more comprehensive representation of beam behaviour [reissner_one-dimensional_1972, reissner_finite_1981, simo_finite_1985]. While numerous finite element formulations [simo_three-dimensional_1986, zupan_quaternion-based_2009, jelenic_geometrically_1999, ibrahimbegovic_computational_1995] have been introduced in the literature using geometrically exact Simo-Reissner beam theory, Tuković et al. [tukovic_finite_2019] were the first to propose a finite volume formulation of geometrically exact Simo-Reissner beams. More recently, Bali et al. [bali_cell-centered_2022] expanded the finite volume formulation for shear-deformable geometrically exact beams and conducted a comparative analysis with finite element-based classical benchmark cases using OpenFOAM. Additionally, they extended this formulation to handle beam-to-beam contact interactions [bali_finite_2024]. The primary motivation for adapting the Simo–Reissner formulation to finite volume discretisation within OpenFOAM is to unify numerical methods for solids and fluids, allowing complex multi-physics problems, like fluid-solid interactions, to be addressed within a unified framework. Furthermore, bali_cell-centered_2022 noted that, unlike the finite element approach, the finite volume formulation for beams is not subject to the shear-locking phenomenon.
@@
In the current work, the finite volume Simo–Reissner beam formulation introduced by bali_cell-centered_2022 is extended by fully integrating it with a multiphase CFD solver and a six-degree-of-freedom motion solver. This results in a unified, fully coupled fluid–structure interaction framework within the finite volume method. The proposed framework is capable of capturing large nonlinear deformations in mooring lines—including bending, shear, and torsion—while simultaneously resolving the free surface dynamics of the fluid. While many existing approaches simplify mooring dynamics using lumped-mass or quasi-static models, the present work introduces a unified, comprehensive simulation environment for moored floating structures. A key novelty of this study lies in the application of a geometrically exact dynamic beam theory in a finite volume context, which allows for consistent coupling with the CFD solver and effective simulation of nonlinear mooring behaviour. In addition, the model incorporates seabed contact and friction effects to capture realistic boundary interactions. The coupled framework is validated and verified against both experimental measurements [wu_experimental_2019] and numerical results from MoorDyn [chen_cfd_2022] for a floating box moored with catenary lines under regular wave conditions, as part of an introductory test campaign for wave energy converters. The rest of the paper is organised as follows. Section 2 provides the details of the mathematical model of the moored floating body problem and the coupling strategy. Section 3 elaborates on the benchmarking of the proposed algorithm. Finally, some conclusions are drawn in Section 4.
@@ 2 Mathematical Model and Numerical Methods
The proposed approach utilises OpenFOAM [weller_tensorial_1998] as the main framework for CFD analysis, while the beam solver, also implemented within the OpenFOAM framework, is used for mooring line dynamics modelling. Consequently, the equation of motion for the floating rigid body and its interaction with mooring lines must be coupled with the CFD model. This section elaborates on the mathematical models that have been used for each component of a moored floating body problem.
@@ 2.1 Free surface flows
The volume of fluid (VOF) formulation based on the surface capturing method with Reynolds Averaging Navier-Stokes (RANS) equations is employed here for solving a two-phase air-water system. The continuity and the conservation of momentum equations are given by:
@@
Physically, Eqn. 5 represents the conservation of mass for one of the phases in a multiphase flow, ensuring that the volume fraction evolves correctly over time. The Multidimensional Universal Limiter with Explicit Solution (MULES) algorithm [damian_extended_2014] is applied here to reduce interface diffusion. This method introduces an artificial compression (or anti-diffusion) term $\nabla\cdot[\mathbf{U_{r}}\alpha(1-\alpha)]$ on the left side of Eqn. 5. In this term, $\mathbf{U_{r}}$ represents the compression velocity. The final expression for the volume fraction advection equation then becomes:
@@
Further details on the MULES formulation can be found in damian_extended_2014.
@@ 2.2 Mooring model
$\frac{\partial\mathbf{m}}{\partial s}\;+\;\mathbf{r}^{\prime}\times\mathbf{n}\;+\;\mathbf{t}\;=\;\mathbf{I}_{\rho}\,\frac{\partial\bm{\omega}}{\partial t}\;+\;\bm{\omega}\times\!\left(\mathbf{I}_{\rho}\,\bm{\omega}\right)$ (8)
where in Eqn. 7, $\mathbf{n}(s,t)$ represents the internal force resultant vector at position $s$ and time $t$, $\mathbf{f}_{ext}(s,t)$ is the external force per unit length acting on the beam at position $s$ and time $t$, $\rho_{b}$ is the density of the beam, and $A$ is the cross-sectional area of the beam. Consequently, $\rho_{b}A$ is the mass per unit length of the beam, $\frac{\partial^{2}\mathbf{r}}{\partial t^{2}}$ is the linear acceleration of the point on the centre line at position $s$. In the moment equation (Eqn. 8), $\mathbf{m}$ is the internal moment vector, $\mathbf{r}^{\prime}$ is the tangent vector along the beam’s centre line, $\bm{\tau}_{\text{ext}}(s,t)$ represents the external torques, $\mathbf{I}_{\rho}$ is the spatial time-dependent inertia tensor of the beam’s cross-section and $\bm{\omega}$ is the angular velocity vector. Equation 7 represents the conservation of linear momentum, where the net internal and external forces acting on a beam segment result in its linear acceleration. Similarly, Eqn. 8 represents the conservation of angular momentum, accounting for the net internal moments and external torques acting on the segment.
@@
The drag force $\mathbf{f}_{d}$ based on Morrison’s equation [morison_force_1950], which describes the total hydrodynamic forces acting on a structure submerged in or interacting with fluid flow, comprises normal and tangent components expressed as [palm_coupled_2016]
@@
[…]
$\mathbf{V}_{r,t}$ and $\mathbf{V}_{r,n}$ are the tangential and normal components of the relative velocity between the beam and the surrounding fluid, i.e. $\mathbf{V}_{r}=\mathbf{V}-\mathbf{V}_{fluid}$. In this study, a quiescent fluid is assumed around the mooring line, i.e. $\mathbf{V}_{fluid}=0$. Similar to the drag force, the added mass force is one of the components in the Morison equation [morison_force_1950]. The added mass force arises because when a body moves through a fluid, it must push aside or accelerate the surrounding fluid. This extra inertia, which the body feels due to the fluid’s resistance, is called added mass. This term increases the effective mass of the beam because it includes both the actual mass of the beam and the mass of the fluid that the beam displaces during motion. Following the similar notation in Eqn. 10, the added mass force can be expressed as
@@
The total buoyancy force acting on the beam, proportional to the displaced fluid, can be expressed as [palm_coupled_2016]
@@
where $\mathbf{A}_{w}$, $\mathbf{A}_{c}$, and $\mathbf{A}_{e}$ are $6\times 6$ coefficient matrices; $\mathbf{X}_{w}$, $\mathbf{X}_{c}$, and $\mathbf{X}_{e}$ are $6\times 1$ vectors of unknown displacements and rotations, and $\mathbf{B}_{c}$ is the corresponding $6\times 1$ source vector. This form is assembled for each beam cell, leading to a global system of $6\times N$ coupled equations. The resulting global matrix is sparse and block tri-diagonal, and the non-linear system is solved using the Newton–Raphson method. The detailed formulation of the Simo-Reissner beam theory, alongside its discretisation, can be found in Bali et al. [bali_cell-centered_2022].